OpenStudy (anonymous):

the sum of the measures of angle x and angle y is 90. If the measure of angle x is 30 less than twice the measure of angle y, what is the measure of angle x?

OpenStudy (anonymous):

so to start you will need to make the algebraic equation for the measure of the angles

OpenStudy (anonymous):

x + y = 90. y in the line 2 with (90 - x) x -30 = 180 - 30

OpenStudy (anonymous):

x+y=90 x=2y-30 then just fill in x in the first equation with 2y-30 and solve it :)